OpenStudy (anonymous):

7/x - 2e^x Find an anti-derivative of f(x) I almost have it correct I think but I'm messing something up. I moved x to the top making it 7x^-1 - 2e^x but I am unsure how take the derivative of -2e^x

OpenStudy (anonymous):

wolframalpha my friend

OpenStudy (anonymous):

http://www.wolframalpha.com/

OpenStudy (savvy):

\[7\ln x - 2e ^{x} + c\]

OpenStudy (anonymous):

thank you! how did you know to to change 7 into a natural log?

OpenStudy (savvy):

dont write 1/x as x^-1..... instead remember the anti derivative of 1/x is \[\ln x\].. and the anti derivative of \[e ^{x}\] is \[e ^{x}\] itself...

OpenStudy (savvy):

i didn't change 7 into natural log....it was 1/x which got transformed to lnx through integration...

OpenStudy (zarkon):

I would write the antiderivative of \(\displaystyle \frac{1}{x}\) as \(\displaystyle\ln|x|+c\)